Operations Report – December 12th

Crew 200 Operations Report 12-Dec-2018

SOL: 10

Name of person filing report: Antoine Bocquier

Non-nominal systems: generator autostart
Notes on non-nominal systems: None

Generator (hours run): From 5 pm last night to 10:40 am (SOC was still
at 78% this morning, not changing from yesterday night). On tonight at

Solar— SOC 72%
Diesel Reading – 100 % (refilled)
Propane Reading – 30 %
Ethanol Free Gasoline – N/A gallons

Water (auxillary tank) – 0 gallons
Water (static tank) – 350 gallons
Auxillary to Static tank transfer – no
Gallons transferred: 0
Water in GreenHab – high level
Water (loft) – Static to Loft Pump used – yes
Water Meter: 3

Toilet tank emptied: no

Deimos rover used: no, still not functional
Hours: 141.5
Beginning charge:
Ending charge: 70%
Currently charging: no

Sojourner rover used: No
Beginning charge:
Ending charge: 100%
Currently charging: yes

Spirit rover used: yes (by media)
Hours: 60.0
Beginning charge: 100%
Ending charge: 54%
Currently charging: yes

Opportunity rover used: yes (by media)
Hours: 41.8
Beginning charge: 100%
Ending charge: 88%
Currently charging: yes

Curiosity rover used: yes (by media)
Hours: not checked
Beginning charge: 100%
Ending charge: not checked
Currently charging: yes

Notes on rovers: Rovers were used today by the media, we could not use
Opportunity for our planned EVA as they were using it. Curiosity was
not present at the base when we came back from EVA, where we checked
the batteries levels.
# Hours the ATVs were Used today: 0
Notes on ATVs: ATVs were taken out of active service and are now
parked away from the station

HabCar used and why, where? No
CrewCar used and why, where? No

General notes and comments: This afternoon and evening are very windy,
we had to securize some of the tunnels tarps with zips. One on the RAM
tunnel had to be removed and placed in the RAM. Temperature in the hab
dropped from 23°C to 19°C, heater is working, it is improving.

Summary of internet: Nothing to report.

Summary of suits and radios: we performed an EVA this morning to
study the radio coverage, everything was alright. Ventilation of
spacesuit 10 sometimes stops before starting later. Spacesuit 3 still
not functional.

Summary of Hab operations: we catched a mouse yesterday night, which
we released today far from the Hab during EVA.

Summary of GreenHab operations: Regular activities

Summary of ScienceDome operations: Nothing to report.

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Musk Solar observatory – Nominal.
Robotic Observatory – Nominal

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report

Copyright © The Mars Society. All rights reserved. | Main Site